OpenStudy (anonymous):

The base of a rectangular tank is 2 feet wide and 4 feet long; the height of the tank is 20 inches. If water is pouring into the tank at the rate of 2 cubic inches per second, how many hours will be required to fill the tank?

OpenStudy (anonymous):

can you find the volume of the tank in cubic inches?

OpenStudy (anonymous):

160 cubed equals 4096000

OpenStudy (anonymous):

two cubic inches is 3456 inches

OpenStudy (anonymous):

the tank is 4096000 in feet, right?

OpenStudy (anonymous):

but in cubic feet it is 160

OpenStudy (anonymous):

wait, I might be confusing myself

OpenStudy (anonymous):



OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

k

OpenStudy (anonymous):

23040

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

2 inch^3/sec * t sec = 23040 inch^3

OpenStudy (anonymous):

t=?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

23040/2=11520/60=192/60=3.2?

OpenStudy (anonymous):

2*t =23040 t=?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

why am I having to divide by 60 twice in order to get the answer?

OpenStudy (anonymous):

we found how many seconds it takes to fill the tank. The question asks how many hours.

OpenStudy (anonymous):

right

OpenStudy (anonymous):

the first I divided by 60, I got 192

OpenStudy (anonymous):

seconds/60 = minutes minutes/60 = hours

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

Now I see

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

sure

OpenStudy (phi):

looks like you got it.

OpenStudy (anonymous):

ty for helping me

OpenStudy (anonymous):

ty